OpenStudy (anonymous):

multiply(x – 8)(x + 12)

OpenStudy (anonymous):

\[x^2+12x-8x-96=x^2+4x-96\]

jhonyy9 (jhonyy9):

x2 -8x+12x-96=x2+4x-96

OpenStudy (anonymous):

(x – 8)(x + 12) \[=x ^{2}+4x-96\]